# On an integrable discretization of the massive Thirring model in light-cone coordinates and the associated Yang–Baxter map

### v1

@@ 1 Introduction
In this paper, we aim to construct a Yang–Baxter map that admits a natural continuous limit. Such a map $R$ is generally supposed to involve an arbitrary (typically small) parameter $\hbar$ and takes the following form [5]:
@@
where $I$ is the identity mapping.
@@
where the subscripts $t$ and $x$ denote the partial differentiation with respect to these variables. More generally, if the map $R$ admits the expansion in terms of (typically small) parameters $h$ and $\varDelta$ as
$R:(q,u)\to\left(\widehat{q},\widehat{u}\right):=(q,u)+\left(hf(q,u),\varDelta g(q,u)\right)+\mathcal{O}(h^{2},h\varDelta,\varDelta^{2}),$ (1.3)
then the map $R$ provides in the continuous limit $h\to 0$ or $\varDelta\to 0$ a semi-discrete system in $1+1$ space-time dimensions.
The asymptotic expansion (1.2) with respect to $\hbar$ is a natural ansatz in that the quantum Yang–Baxter equation (1.1) for a linear operator $R$ reduces in the quasi-classical limit $\hbar\to 0$ to the classical Yang–Baxter equation [6, 7]:
$[r_{12},r_{13}]+[r_{12},r_{23}]+[r_{13},r_{23}]=0,$ (1.4)
where [,][\;\,,\;\,][ , ] is the commutator.
In our previous paper [8], we already constructed a Yang–Baxter map of the form (1.3) (see also some relevant results in [9, 10, 11, 12, 13, 14]). In this paper, we propose another Yang–Baxter map of the form (1.3), which appears to be more general and interesting than the Yang–Baxter map constructed in our previous paper [8]; closely related results were obtained previously by Xu and Pelinovsky [15], but they did not explicitly present their results in the form of a Yang–Baxter map.
@@ 4 Yang–Baxter map
and consider the simplest case of the scalar dependent variables, we obtain the Yang–Baxter map $\left(q_{n},u_{n+1}\right)\mapsto\left(\widetilde{q}_{n},u_{n}\right)$ in the form:
@@
Clearly, (4.30) can be interpreted as a fully discrete analog of the massive Thirring model in light-cone coordinates (2.16), so the Yang–Baxter map $\left(q_{n},u_{n+1}\right)\mapsto\left(\widetilde{q}_{n},u_{n}\right)$ indeed belongs to the class discussed in the introduction.
@@ 5 Concluding remarks
In this paper, we use a discrete Lax matrix representing the binary Bäcklund–Darboux transformation for the continuous derivative nonlinear Schrödinger (Chen–Lee–Liu[19]) hierarchy as the main building block. The corresponding discrete zero-curvature condition can be solved to provide a fully discrete analog of the massive Thirring model in light-cone (or characteristic) coordinates. This discrete massive Thirring model indeed reduces to the continuous massive Thirring model in a suitable continuous limit. The binary Bäcklund–Darboux transformation for the discrete massive Thirring model can be constructed in a straightforward manner and the one-soliton solution is presented explicitly. Using some symmetry properties of the discrete Lax matrix, we can rewrite the discrete massive Thirring model as a parameter-dependent Yang–Baxter map. The Yang–Baxter map satisfies the parameter-dependent version [36, 10] of the quantum Yang–Baxter equation (1.1), i.e.,
@@
Thus, we obtain a new parameter-dependent Yang–Baxter map that admits a natural continuous limit; the two parameters therein can be understood as the lattice parameters. Note that this Yang–Baxter map can also be constructed using the definition of the unknown functions appearing in the discrete Lax matrix in terms of the linear eigenfunction (see (3.1) and (3.12) with (3.19)). It would be interesting to discuss how this Yang–Baxter map can be characterized within the framework of the matrix modified KP equation (see [11, 12] for the simpler case of the matrix KP equation).

### v2

@@ 1 Introduction
In this paper, we aim to construct a Yang–Baxter map that admits a natural continuous limit. In the simplest case, such a map $R$ involves an arbitrary (typically small) parameter $\hbar$ and takes the following form [5]:
@@
where $I$ is the identity mapping. The asymptotic expansion (1.2) with respect to $\hbar$ is a natural ansatz in that the quantum Yang–Baxter equation (1.1) for a linear operator $R$ reduces in the quasi-classical limit $\hbar\to 0$ to the classical Yang–Baxter equation [6, 7]:
$[r_{12},r_{13}]+[r_{12},r_{23}]+[r_{13},r_{23}]=0,$
where [,][\;\,,\;\,][ , ] is the commutator.
@@
where the subscripts $t$ and $x$ denote the partial differentiation with respect to these variables. In our previous paper [8], we already constructed a Yang–Baxter map of this type (see also some relevant results in [9, 10, 11, 12, 13, 14]).
We can consider a more general situation where the map $R_{12}$ admits the expansion in terms of (typically small) parameters $h$ and $\varDelta$ as
$R_{12}:(q,u)\to\left(\widehat{q},\widehat{u}\right):=(q,u)+\left(hf(q,u),\varDelta g(q,u)\right)+(\mathcal{O}(h^{2},h\varDelta),\mathcal{O}(h\varDelta,\varDelta^{2})).$ (1.3)
In this paper, we propose a Yang–Baxter map that admits the expansion (1.3) and provides in the continuous limit $h\to 0$ or/and $\varDelta\to 0$ a nontrivial semi-discrete/continuous system in $1+1$ space-time dimensions. The Yang–Baxter map proposed in this paper appears to be more general and interesting than the Yang–Baxter map constructed in our previous paper [8]; closely related results were obtained previously by Xu and Pelinovsky [15], but they did not explicitly present their results in the form of a Yang–Baxter map.
@@ 4 Yang–Baxter map
and consider the simplest case of the scalar dependent variables, we obtain the map $\left(q_{n},u_{n+1}\right)\mapsto\left(\widetilde{q}_{n},u_{n}\right)$ in the form:
@@
Clearly, (4.30) can be interpreted as a fully discrete analog of the massive Thirring model in light-cone coordinates (2.16), so the Yang–Baxter map $\left(q_{n},u_{n+1}\right)\mapsto\left(\widetilde{q}_{n},u_{n}\right)$ given by (4.27) and (4.28) indeed belongs to the class discussed in the introduction.
@@ 5 Concluding remarks
In this paper, we use a discrete Lax matrix representing the binary Bäcklund–Darboux transformation for the continuous derivative nonlinear Schrödinger (Chen–Lee–Liu[19]) hierarchy as the main building block. By solving the corresponding discrete zero-curvature condition, we obtain a fully discrete system, which provides, with a particular choice of the parameters, an integrable discretization of the massive Thirring model in light-cone (or characteristic) coordinates. The binary Bäcklund–Darboux transformation for the discrete massive Thirring model can be constructed in a straightforward manner and the one-soliton solution is presented explicitly. Using some symmetry properties of the discrete Lax matrix, we can rewrite the fully discrete system, which includes the discrete massive Thirring model as a particular case, as a parameter-dependent Yang–Baxter map. The Yang–Baxter map satisfies the parameter-dependent version [36, 10] of the quantum Yang–Baxter equation (1.1), i.e.,
@@
Thus, we obtain a new parameter-dependent Yang–Baxter map that admits a natural continuous limit for $R_{12}$; the two parameters therein can be understood as the lattice parameters. Note that this Yang–Baxter map can also be constructed using the definition of the unknown functions appearing in the discrete Lax matrix in terms of the linear eigenfunction (see (3.1) and (3.12) with (3.19)). It would be interesting to discuss how this Yang–Baxter map can be characterized within the framework of the matrix modified KP equation (see [11, 12] for the simpler case of the matrix KP equation).